Compute the dosage in rems if a person receives 12 rads of

a) X-rays

b) γ rays

c) slow neutrons (RBE = 2)

d) fast neutrons to the body

e) fast neutrons to the eyes
Answer #1

Radiation dose equivalent = absorbed dose * RBE factor

absorbed dose = 12 rads

a), b) x rays and gama rays RBE factor =1

dose equivalent = 12 *1 = 12 Rem

c) slow neutrons (< 1Mev) RBE = 2.5+ 18.2exp(-[lnE]2 /6)

= 2.5 + 18.2 exp( -ln12/6) = 20.7

dose equivalent = 12*20.7 = 248.4 Rem

d) fast neutrons whole body , assume E = 10 Mev

RBE = 5.0 + 17.0 exp (-ln20)2/6 ) = 8.81

dose equivalent H = 12*8.81 = 105.72 Rem

e) fast neutrons to the eyes

Tissue weight factor for eyes = 0.12

Dose equivalent = 105.71 *0.12 = 12.69 Rem